# Reinforcement Learning meets Masked Video Modeling : Trajectory-Guided Adaptive Token Selection

### v1

@@ Abstract
Figure 1: A depicts our overall architecture with MAE ($f_{\phi}$) and TATS ($g_{\theta}$). B illustrates the joint training (Algorithm 1) of $f_{\phi}$ and $g_{\theta}$ using PPO. Until epoch $m_{o}$, standard random space-time masking is applied. Afterward, every $k$ steps, Phase 1 ($g_{\theta}$ frozen,$f_{\phi}$ unfrozen) stores old state of $g_{\theta}$ in memory buffer $\mathcal{M}_{b}$ as episodes, followed by Phase 2 ($g_{\theta}$ unfrozen,$f_{\phi}$ frozen), where $g_{\theta}$ is optimized via $\mathcal{L}_{s}(\theta)$. The optimization process then alternates between Phase 1 and Phase 2.
@@ 3.1 Overview of MVM
Tokenizer. Consider an input video $V$ of size $T\times C\times H\times W$, where $T$ represents the number of frames, $C$ denotes the input channels and $H$, $W$ is the height and width of a frame. A Tokenizer composed of 3D convolutional layer with kernel $K$ of size $(t,C,h,w)$, stride $S$ of size $(t,h,w)$ and $d$ output channels is availed to tokenize $V$ into $N$ number of tokens with dimensiion $d$ indicated as $\mathbf{X}$, where $N=\frac{T}{t}\times\frac{H}{h}\times\frac{W}{w}$. Positional information is further embedded into the tokens using a fixed 3D periodic positional encoding scheme outlined in [54].
@@ 3.2 Trajectory-Aware Adaptive Token Sampler
Here, $\exp$ denotes the exponential function, .,$\langle$.,$\rangle$ . , . ⟩ represents dot product, and $\cdot$ indicates element-wise multiplication. Next, trajectories are pooled across time to capture intra-frame relationships. The trajectory tokens $\tilde{\mathbf{y}}_{stt^{\prime}}$ are mapped to a new set of query, key, and value representations, denoted as $\tilde{\mathbf{q}}_{st},\tilde{\mathbf{k}}_{stt^{\prime}},\tilde{\mathbf{v}}_{stt^{\prime}}$, using the projection matrix $\tilde{W}$. Now $\tilde{\mathbf{q}}_{st}$ becomes the updated reference query for reference point $st$. $\tilde{\mathbf{q}}_{st}$ is then utilized to aggregate information across the temporal dimension using 1D attention given by:
@@
$\boldsymbol{Z}=\text{TA}(\boldsymbol{X})\text{; }\text{Z}\in\mathbb{R}^{N\times d}\$ (3)
@@ 3.3 Optimization
$\mathcal{L}_{R}(\phi)=\frac{1}{N-N_{v}}\sum_{i\in\boldsymbol{I}_{m}}||\mathbf{\tilde{V}}_{i}-\mathbf{\hat{V}}_{i}||_{2}$ (6)
@@
1:Video $V$, MAE network $f_{\phi}$, TATS module $g_{\theta}$, mask ratio $\rho$, memory buffer $\mathcal{M}_{b}$, epochs $E$, Train only MAE epochs $m_{o}$, TATS update interval $k$, Total number of tokens $N$.
@@ 4 Experimental Setup
Implementation Details. We employ the ViT-Base model ($\approx$87M parameters) [11] for our experiments. The input video has the dimension $16\times 3\times 224\times 224$ while the patch size is $2\times 3\times 16\times 16$ (tubelet length $=2$), yielding a total of $1568$ tokens. Mask ratio $\rho$ takes the value $\{0.85,0.90,0.95\}$. Our experiments contain two types of settings:
@@
Refer to the supplementary material (supp) for additional implementation details.
Table 1: Comparison of fine-tuning result of Our model against baselines ([4, 52]) on action recognition task across benchmark datasets and different $\rho$ with top-1/top-5 accuracy as evaluation metric. (($\uparrow$ ) / ($\downarrow$ ) : denotes increase/decrease in performance)
. Dataset Mask Ratio VideoMAE [52] AdaMAE [4] Ours $\rho$ top-$1$ top-$5$ top-$1$ top-$5$ top-$1$ top-$5$ 0.85 80.36 94.95 83.98 96.37 85.94 ($\uparrow$ ) 96.98 ($\uparrow$ ) UCF101 0.90 76.64 94.29 82.42 95.84 84.53 ($\uparrow$ ) 96.37 ($\uparrow$ ) 0.95 65.86 89.14 80.83 95.26 81.75 ($\uparrow$ ) 95.29 ($\uparrow$ ) 0.85 40.82 71.61 41.28 73.37 41.60 ($\uparrow$ ) 73.31 ($\downarrow$ ) HMDB51 0.90 36.39 69.73 39.13 72.33 41.28 ($\uparrow$ ) 73.76 ($\uparrow$ ) 0.95 33.98 65.36 37.70 70.38 38.67 ($\uparrow$ ) 72.01 ($\uparrow$ ) 0.85 42.26 68.28 38.97 64.68 43.24 ($\uparrow$ ) 68.76 ($\uparrow$ ) Kinetics-400 0.90 41.79 68.62 39.50 65.70 43.28 ($\uparrow$ ) 68.85 ($\uparrow$ ) 0.95 39.73 66.15 39.42 65.14 41.70 ($\uparrow$ ) 67.29 ($\uparrow$ ) 0.85 37.63 66.47 37.92 66.63 39.96 ($\uparrow$ ) 68.10 ($\uparrow$ ) SSv2 0.90 37.85 66.86 38.10 66.29 40.79 ($\uparrow$ ) 69.30 ($\uparrow$ ) 0.95 37.24 65.92 38.38 67.11 40.25 ($\uparrow$ ) 68.73 ($\uparrow$ )
@@ 4.2 Ablation Studies
Table 2: Comparison of transfer learning result of Our model against [4, 52] on action recognition across benchmark datasets and different $\rho$ with top-1/top-5 accuracy as evaluation metric. (($\uparrow$ ) / ($\downarrow$ ) / (-- ) : denotes increased/decreased/equivalent performance)
@@
[…]
@@
4. Effect of Number of Trajectory Attention Blocks. In Table LABEL:tab:TATS_network, we investigate the effect of varying the # of TA blocks in TATS. Our results indicate that the configuration with # TA Blocks $=1$ yields the highest top-1 accuracy of $81.75\%$. As we increase the # TA Blocks, the performance decreases while the memory usage increases.
@@ 7 Datasets
Learning rate for $g_{\theta}$ $lp$ 1.5e-6
@@ 8.2 Hyper-parameter Setting
Pre-training. The hyperparameter configurations used during the pre-training phase across all benchmark datasets are presented in Table 4. For $(m_{o},k)$, hyperparameter tuning is conducted on the UCF101 and HMDB51 datasets (Table 5), and the configuration that minimizes the reconstruction error is selected. Similarly we also perform hyperparameter tuning for coefficients ($c_{1}$,$c_{2}$,$c_{3}$) in Table 6 during pretraining on UCF101 and observe that (1e-4, 1e-4, 1e-4) minimizes the reconstruction error. Empirical observations indicate that the optimal configuration for UCF101 also performs effectively on subset of K400 and SSv2 (small scale pre-training setup). It is to be noted that we use reconstruction loss for tuning these hyper-parameters because behaviour of reconstruction loss during pretraining is more interpretable in terms of convergence than the sampling loss.
@@
Input Video stride ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4}\times{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}1}\times{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}1}$ for K400 ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3}\times{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}16}\times{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}224}\times{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}224}$
stride ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2}\times{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}1}\times{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}1}$ for ssv2/ucf/hmdb
@@
Masking TATS Masking $[(1-\rho)\times{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}1568}]\times{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}768}$
@@
Encoder $[MHA({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}768})]\times 12$ $[(1-\rho)\times{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}1568}]\times{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}768}$
@@
Decoder $[MHA({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}384})]\times 4$ $[(1-\rho)\times{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}1568}]\times{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}384}$
@@
Reshaping from 1536 to ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3}\times{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2}\times{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}16}\times{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}16}$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}3}\times{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}16}\times{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}224}\times{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}224}$
@@ 9 Large Scale Pre-training Results
Table 9: Large Scale Pre-training and Finetuning Results. Comparison of fine-tuning result of Our model against baselines ([4, 52]) on action recognition task for full SSv2 and $\rho=0.95$ with top-1/top-5 accuracy as evaluation metric. (($\uparrow$ ) : denotes increase in performance)

### v2

@@ Abstract
[…]
@@ 3.1 Overview of MVM
Tokenizer. Consider an input video $V$ of size $T\times C\times H\times W$, where $T$ represents the number of frames, $C$ denotes the input channels and $H$, $W$ is the height and width of a frame. A Tokenizer composed of 3D convolutional layer with kernel $K$ of size $(t,C,h,w)$, stride $S$ of size $(t,h,w)$ and $d$ output channels is availed to tokenize $V$ into $N$ number of tokens with dimension $d$ indicated as $\mathbf{X}$, where $N=\frac{T}{t}\times\frac{H}{h}\times\frac{W}{w}$. Positional information is further embedded into the tokens using a fixed 3D periodic positional encoding scheme outlined in [54].
@@ 3.2 Trajectory-Aware Adaptive Token Sampler
Here, $\exp$ denotes the exponential function, $\langle.,.\rangle$ represents dot product, and $\cdot$ indicates element-wise multiplication. Next, trajectories are pooled across time to capture intra-frame relationships. The trajectory tokens $\tilde{\mathbf{y}}_{stt^{\prime}}$ are mapped to a new set of query, key, and value representations, denoted as $\tilde{\mathbf{q}}_{st},\tilde{\mathbf{k}}_{stt^{\prime}},\tilde{\mathbf{v}}_{stt^{\prime}}$, using the projection matrix $\tilde{W}$. Now $\tilde{\mathbf{q}}_{st}$ becomes the updated reference query for reference point $st$. $\tilde{\mathbf{q}}_{st}$ is then utilized to aggregate information across the temporal dimension using 1D attention given by:
@@
$\boldsymbol{Z}=\text{TA}(\boldsymbol{X})\text{; }\text{Z}\in\mathbb{R}^{N\times d}\\$ (3)
@@ 3.3 Optimization
[…]
@@
1:Video $V$, MAE network $f_{\phi}$, TATS module $g_{\theta}$, mask ratio $\rho$, memory buffer $\mathcal{M}_{b}$, epochs $E$, Train only MAE epochs $m_{o}$, TATS update interval $k$, Total number of tokens NN.
@@ 4 Experimental Setup
Implementation Details. We employ the ViT-Base model ($\approx$87M parameters) [11] for our experiments. The input video has the dimension $16\times 3\times 224\times 224$ while the patch size is $2\times 3\times 16\times 16$ (tubelet length =2=2), yielding a total of $1568$ tokens. Mask ratio $\rho$ takes the value $\{0.85,0.90,0.95\}$. Our experiments contain two types of settings:
@@
Refer to the supplementary material (supp) for additional implementation details. Source code for this work has been released at github.com/rayush7/rl_videomae.
Table 1: Comparison of fine-tuning result of Our model against baselines ([4, 52]) on action recognition task across benchmark datasets and different $\rho$ with top-1/top-5 accuracy as evaluation metric. $(\uparrow$ ) / ($\downarrow$ ) : denotes increase/decrease in performance)
. Dataset Mask Ratio VideoMAE [52] AdaMAE [4] Ours $\rho$ top-11 top-55 top-11 top-55 top-11 top-55 0.85 80.36 94.95 83.98 96.37 85.94 ($\uparrow$ ) 96.98 ($\uparrow$ ) UCF101 0.90 76.64 94.29 82.42 95.84 84.53 ($\uparrow$ ) 96.37 ($\uparrow$ ) 0.95 65.86 89.14 80.83 95.26 81.75 ($\uparrow$ ) 95.29 ($\uparrow$ ) 0.85 40.82 71.61 41.28 73.37 41.60 ($\uparrow$ ) 73.31 ($\downarrow$ ) HMDB51 0.90 36.39 69.73 39.13 72.33 41.28 ($\uparrow$ ) 73.76 ($\uparrow$ ) 0.95 33.98 65.36 37.70 70.38 38.67 ($\uparrow$ ) 72.01 ($\uparrow$ ) 0.85 42.26 68.28 38.97 64.68 43.24 ($\uparrow$ ) 68.76 ($\uparrow$ ) Kinetics-400 0.90 41.79 68.62 39.50 65.70 43.28 ($\uparrow$ ) 68.85 ($\uparrow$ ) 0.95 39.73 66.15 39.42 65.14 41.70 ($\uparrow$ ) 67.29 ($\uparrow$ ) 0.85 37.63 66.47 37.92 66.63 39.96 ($\uparrow$ ) 68.10 ($\uparrow$ ) SSv2 0.90 37.85 66.86 38.10 66.29 40.79 ($\uparrow$ ) 69.30 ($\uparrow$ ) 0.95 37.24 65.92 38.38 67.11 40.25 ($\uparrow$ ) 68.73 ($\uparrow$ )
@@ 4.2 Ablation Studies
Table 2: Comparison of transfer learning result of Our model against [4, 52] on action recognition across benchmark datasets and different $\rho$ with top-1/top-5 accuracy as evaluation metric. $(\uparrow$ ) / ($\downarrow$ ) / (- ) : denotes increased/decreased/equivalent performance)
@@
TA (# Blocks=1)($\text${\# Blocks}=1) 81.75 95.29 25.36 GB
TA (# Blocks=2)($\text${\# Blocks}=2) 65.17 88.59 32.18 GB
TA (# Blocks=3)($\text${\# Blocks}=3) 67.35 90.20 39.00 GB
@@
4. Effect of Number of Trajectory Attention Blocks. In Table LABEL:tab:TATS_network, we investigate the effect of varying the # of TA blocks in TATS. Our results indicate that the configuration with # TA Blocks =1=1 yields the highest top-1 accuracy of $81.75\%$. As we increase the # TA Blocks, the performance decreases while the memory usage increases.
@@ 7 Datasets
Learning rate for $g_{\theta}$ - lplp 1.5e-6
@@ 8.2 Hyper-parameter Setting
Pre-training. The hyperparameter configurations used during the pre-training phase across all benchmark datasets are presented in Table 4. For $(m_{o},k)$, hyperparameter tuning is conducted on the UCF101 and HMDB51 datasets (Table 5), and the configuration that minimizes the reconstruction error is selected. Similarly we also perform hyperparameter tuning for coefficients ($c_{1}$$c_{2}$$c_{3}$) in Table 6 during pretraining on UCF101 and observe that (1e-4, 1e-4, 1e-4) minimizes the reconstruction error. Empirical observations indicate that the optimal configuration for UCF101 also performs effectively on subset of K400 and SSv2 (small scale pre-training setup). It is to be noted that we use reconstruction loss for tuning these hyper-parameters because behaviour of reconstruction loss during pretraining is more interpretable in terms of convergence than the sampling loss.
@@
[…]
@@
Masking TATS Masking $(1-\rho)\times{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}1568}]\times{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}768}$
@@
Encoder $MHA({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}768})]\times 12 [$$(1-\rho)\times{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}1568}]\times{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}768}$
@@
Decoder $MHA({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}384})]\times 4 [$$(1-\rho)\times{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}1568}]\times{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}384}$
@@
[…]
@@ 9 Large Scale Pre-training Results
Table 9: Large Scale Pre-training and Finetuning Results. Comparison of fine-tuning result of Our model against baselines ([4, 52]) on action recognition task for full SSv2 and $\rho=0.95$ with top-1/top-5 accuracy as evaluation metric. $(\uparrow$ ) : denotes increase in performance)